# Link Prediction with Physics-Inspired Graph Neural Networks thanks:

### v2

@@ Abstract
The message-passing mechanism underlying Graph Neural Networks (GNNs) is not naturally suited for heterophilic datasets, where adjacent nodes often have different labels. Most solutions to this problem remain confined to the task of node classification. In this article, we focus on the valuable task of link prediction under heterophily, an interesting problem for recommendation systems, social network analysis, and other applications. GNNs like GRAFF have improved node classification under heterophily by incorporating physics biases in the architecture. Similarly, we propose GRAFF-LP, an extension of GRAFF for link prediction. We show that GRAFF-LP effectively discriminates existing from non-existing edges by learning implicitly to separate the edge gradients. Based on this information, we propose a new readout function inspired by physics. Remarkably, this new function not only enhances the performance of GRAFF-LP but also improves that of other baseline models, leading us to reconsider how every link prediction experiment has been conducted so far. Finally, we provide evidence that even simple GNNs did not experience greater difficulty in predicting heterophilic links compared to homophilic ones. This leads us to believe in the necessity for heterophily measures specifically tailored for link prediction, distinct from those used in node classification. The code for reproducing our experiments is available at this URL https://anonymous.4open.science/r/Link_Prediction_with_PIGNN_IJCNN-F03F/.
@@ I Introduction
Most GNNs rely on the message-passing formalism [1]. In heterophilic graphs, where connected nodes tend to have different labels, Message-Passing Neural Networks (MPNNs) may struggle in the classification task as they tend to generate similar representations for adjacent nodes [2], an issue commonly known as over-smoothing [3]. There have been efforts aimed at improving the performance of GNNs on heterophilic graphs [2, 4, 5, 6], however, these works are limited to the context of node classification. An example of this is the gradient flow framework (GRAFF) [7], which deals with heterophily through physics-inspired biases. Although GRAFF shows competitive performance in both heterophilic and homophilic graphs, existing research has focused solely on node classification tasks, not investigating its potential in link prediction— which has significant applications in several domains. This task becomes particularly complex for heterophilic graphs, where we may have a link between two nodes with dissimilar characteristics. Unlike homophilic graphs where connected nodes are similar, the reason behind the connection of two dissimilar entities can be latent. To the best of our knowledge, link prediction under heterophily has only been discussed and brought to the community’s attention by [8].
@@
    We propose a novel Physics-Inspired readout function, that leads to consistent performance improvements for GRAFF-LP as well as other baseline models. Additionally, the new readout gives GRAFF-LP more transparency in its behavior at inference time.
@@ II Related Works
Physics-Inspired vs. Physics-Informed. Physics-Informed (PI) neural networks incorporate physical priors to improve performance and generalizability [14, 15, 16]. Physics-Inspired (PIrd) networks are a subset of PI methods, where physical constraints are embedded in the architecture itself, acting as inductive biases [17, 18, 7, 19].
Physics-Inspired Graph Neural Networks. PIrd GNNs incorporate physics principles directly into the model’s structure. Examples include GNNs based on gradient flows [7], reaction-diffusion equations [20, 21, 19], based on nonlinear controlled and damped oscillators [22], and non-dissipative systems using antisymmetric weight matrices [23]. While these approaches have been applied to node classification, they remain unexplored for link prediction. Our work provides the first perspective on PIrd biases in this setting. The extended discussion on related works is available in Appendix VI.
@@ III Preliminaries
Homophily measures. The homophily assumption in graphs refers to the tendency of similar nodes to be connected. An unambiguous similarity measure is missing in the literature. Those that have been used the most with GNNs are edge homophily $\xi_{edge}$ [24] and node homophily $\xi_{node}$ [2]. The former was more considered within the node classification benchmarks [25, 26, 6, 7], and can be computed as
@@
Both $\xi_{edge}$, and $\xi_{node}$, rely on the labels associated with the nodes. As an example, Equation (1) measures the fraction of edges that connect nodes from the same class. Generally, if we record a low homophily, we consider the graph as heterophilic. Traditional homophily measures are unsuitable for cross-datasets comparison because they are sensitive to class numbers and sample balance [27, 9]. To address this, the adjusted homophily metric $\xi_{adj}$ was introduced in [27]:
@@
Graph Neural Networks as gradient flow
Let us consider an $N$-dimensional dynamic system evolving as $\dot{\mathbf{H}}(t)=F(\mathbf{H}(t))$, with $\mathbf{H}(t)\in\mathbb{R}^{N\times d}$ If there exists a function $E:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}$, s.t. $F(\mathbf{H}(t))=-\nabla E(\mathbf{H}(t))$, the evolution equation $\dot{\mathbf{H}}(t)$ is the gradient flow of the energy $E$. Gradient flows are useful for studying the underlying dynamics of the system, provided the knowledge of $E$. In our case, $N$ represents the graph nodes whose representations $\mathbf{H}(t)$ evolve through a GNN over time, and $E(\mathbf{H}(t))$, is an energy functional associated with the node representations. Let $GNN:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ be an intermediate layer of a generic GNN. By treating the GNN layers as continuous time $t$ and defining $GNN(\mathbf{H}(t))=-\nabla E(\mathbf{H}(t))$, the evolution of the features through the GNN is described as the gradient flow of $E(\mathbf{H}(t))$. $E$ can be selected as the Dirichlet Energy $E^{dir}$:
@@
where $(\nabla\mathbf{H}(t))_{ij}=\frac{\mathbf{h}^{t}_{j}}{\sqrt{D_{jj}+1}}-\frac{\mathbf{h}^{t}_{i}}{\sqrt{D_{ii}+1}}$ and $\mathbf{h}^{t}_{i}$ denotes the feature of node $i$ at time $t$. Poor performance on heterophilic graphs and over-smoothing are often linked to the Dirichlet energy of features decaying to zero as layers increase [28, 29]. Equation (3) shows that decreasing $E^{dir}$ brings adjacent nodes closer in feature space. Conversely from what was commonly thought, [7] proved that linear graph convolutions with symmetric weights shared among layers can induce edge-wise attraction (repulsion) through their positive (negative) eigenvalues. This control mechanism effectively influences whether the features are smoothed or sharpened, making the model successfully handle node classification within heterophilic graphs. This was possible by defining layers as a gradient flow of a parametrized Dirichlet energy:
@@
$\Theta_{+}$ and $\Theta_{-}$ depend on the positive and negative eigenvalues of the weight matrices respectively. By applying Euler discretization to the gradient flow and replacing $\mathbf{H}(t)$ by $\mathbf{H}^{t}$ :
@@
$\tau$ is the integration step, $T=\tau L$ is the total integration time, $\mathbf{\Omega}$, $\mathbf{W}$ and $\tilde{\mathbf{W}}$ are the trainable matrices, that are symmetric and shared across the layers, and $\boldsymbol{\mathbb{A}}=\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{\tilde{A}}\mathbf{\tilde{D}}^{-\frac{1}{2}}$, since self-loops are included in $\mathbf{A}$ and $\mathbf{D}$ (i.e. $\tilde{\mathbf{A}}=\mathbf{I}+\mathbf{A}$). Equation (5), takes the form of a residual network [30], where the interpretation of GNNs as gradient flows is referred to in the literature as PIrd GNN. For details on these derivations, see [7]. In this work, we take advantage of these results and build upon this architecture for link prediction.
@@ IV Proposed Framework: GRAFF-LP
The Hadamard readout is commonly used with node-based link predictors [10]. In this paper, we are the first to propose an alternative function $f_{g}$ which we believe is more compatible with the GRAFF backbone. Since as seen in Equation (4) gradient flow operates by minimizing or maximizing the squared norm of edge gradients and $f_{g}$ can be directly related to them through the following relation:
@@
More precisely, the gradient flow operates by minimizing or maximizing the squared norm of the gradients multiplied by the terms $\Theta_{+}$ and $\Theta_{-}$, we preferred to define the readout function without including those terms to speed up the calculation of $f_{g}$. We use the Hadamard product of the gradients, in order to not provide any bias on the edge directionality since we deal with undirected graphs. However, for directed graphs, we could simply have $f_{g}(\mathbf{z}_{i},\mathbf{z}_{j})=(\nabla\mathbf{H}^{T})_{i,j}$. We do not consider other readout such as concatenation of $z_{i}$ and $z_{j}$, since this would provide a bias on directionality as well, we would lose the physics-inspired bias, and we would double $d_{MLP}$ affecting negatively the space and time complexity.
@@ IV-A Complexity Analysis of GRAFF-LP
We now consider the space complexity in terms of the number of parameters. GRAFF-LP, due to its weight-sharing mechanism, maintains constant complexity with respect to $L$ and exhibits quadratic complexity with respect to the number of hidden dimensions, as the elements in $\mathbf{\Omega}$ and $\mathbf{W}$ are of size $d_{h}^{2}$, where $d_{h}$ is the hidden dimension. Since these matrices are symmetric, there is redundancy in the parameters, reducing the total number to approximately $\frac{1}{2}d_{h}^{2}$.
To summarize, the number of parameters for GCN scale as $\mathcal{O}(Ld_{h}^{2})$. Since GCN shares a similar message-passing as Equation (5), but it does not use weight sharing, and we set $\mathbf{\Omega}\equiv 0$. In the case of GRAFF-LP, both $\mathbf{W}$ and $\mathbf{\Omega}$ use weight sharing, and the overall complexity scales as $\mathcal{O}(d_{h}^{2})$. In terms of time complexity, GRAFF-LP does not have any specific advantage over other models. In Table V, we show the runtime analysis for all the models, reporting also the number of parameters.
@@ V-A Experimental Set-up
Datasets
Table I presents graph statistics after conversion to undirected graphs, a standard GNN procedure. The datasets selected belong to a recent collection [9], which was proposed to enrich the current dataset availability for the GNN experimental setting under heterophily. They have not yet been applied to link prediction. We used four datasets, Amazon Ratings, Roman Empire, Minesweeper, Questions and Tolokers, additional details and dataset descriptions can be found in Appendix II. These datasets differ in context, size, and structural properties [9].
@@
GAT: An attentional MPNN [32] that uses an attention mechanism [33] for neighbor aggregation, particularly effective for heterophilic graphs.
@@
NCNC: [35] A link prediction method, based on structural features as ELPH but these are related to higher order common neighbor information. This method is currently state-of-the-art in a recent link prediction benchmark [13].
@@
Further experimental details are available in Appendix III.
@@ V-B Results
To evaluate model performance, we use the Area Under the Receiver Operating Characteristic (AUROC). Table II summarizes the results across all datasets for both readout functions, $f_{h}$ and $f_{g}$. The results with $f_{h}$ are related to the best configuration of hyperparameters with $f_{h}$. Concerning $f_{g}$, we just took the same configuration found with $f_{h}$ and trained again the models via $f_{g}$. Performance metrics are averaged over 10 random seeds. We adopted a single split since the graphs are sufficiently large and less sensitive to high variance in the data. The statistical significance of the results is computed through the Wilcoxon test [36], via the accuracy of positive and negative edges in the test set, details about the procedure can be found in the code.
@@
With $f_{h}$, GRAFF-LP leads in 3 out of 4 datasets and in 2 out of 4 using $f_{g}$, ranking consistently in the top 2 across all datasets, indicating adaptability not seen with the other models. These aspects applies also for Tolokers, whose results and comments are reported in the Appendix IV-A.
@@
On the Roman Empire dataset, which resembles a chain-like graph, GRAFF-LP achieves the highest AUROC in both setups, with a neat advantage w.r.t. other baselines. ELPH, and NCNC also present a wide gap w.r.t. the node-based methods but they have access to the subgraph features that are a clear advantage to perform link prediction in such structure. This advantage and superiority of ELPH and NCNC are found with Minesweeper, where the graph is a regular grid, and link prediction can be easily solved by understanding the grid structure.
@@
In both cases, the physics-inspired bias seems to let the model use the structure properly and achieve state-of-the-art performance. According to our hyperparameter optimization, ELPH fails to obtain competitive performance in Amazon Ratings and Questions. We relate this gap to the limited depth of ELPH (which is up to 3 layers). NCNC instead, is able to reach competitive performance in Amazon Ratings, as well as Questions. Despite this, it never surpasses GRAFF-LP. We conclude that subgraph-based methods do not perform better than GRAFF-LP on class heterophilic datasets. This implies that structural features are not critical or determining to perform well in these settings.
@@
Another observation is that GCN both in Amazon Ratings and Questions is among the top-scoring models, and in particular gets the closest to GRAFF-LP. This is not unexpected, since they share a similar message-passing and expressivity [37, 7]. They mainly differ in the weight design. We associate these similarities to the same phenomenon observed by [7], who observed that in homophilic node classification, the two models perform the same. We conjecture in this case that for Amazon Ratings and Questions the edge-wise repulsion is not required, thus GCN can afford similar results. We also report the results of the MLP. The MLP results consistently lag behind the GNN models, emphasizing the need for graph representation learning for effective link prediction in these new datasets. For example, in Minesweeper, MLP’s struggles to infer missing links due to its lack of graph context and the dataset’s grid structure.
@@
In Appendix IV-B, we provide examples of GRAFF-LP performing on link prediction under homophily, to show that our approach can reach competitive as well as state-of-the-art performance also under homophily.
In Table III we show the relative percentage improvement we obtained using $f_{g}$ as the readout method. We did not include ELPH and NCNC, since their official derivation is based on Hadamard or other readouts.
@@
GRAFF-LP +0.7% +1.02% 0% 0%
Physics-Inspired Link Prediction
[7] proposed the GRAFF architecture and showed the attraction and repulsion behavior among adjacent nodes with different labels. This is expressed in Equation (4). In the link prediction scenario, we believe that the same behavior can be induced among two nodes that even though they seem different, should attract themselves, and vice versa, where two nodes are apparently similar, but do not present any interaction. Since we build upon the same framework of [7], we expect the edge gradients associated with $\Theta_{+}$ to become smaller, because of attraction in the feature space, and the edge gradients associated with $\Theta_{-}$ to become larger in the feature space because of repulsion. In this way, we can deduce that GRAFF-LP can induce attraction and repulsion in the same fashion as in GRAFF for node classification, and we would answer positively to RQ2. To understand whether this phenomenon happens, we introduce a novel metric, namely the gradient separability ($GS$). Let us assume we have a set of edge gradients =$\nabla=\{(\nabla\mathbf{H}^{t})_{i,j}|0\leq t\leq T\wedge(i,j)\in\mathcal{E}_{t}\}$, where $\mathcal{E}_{t}=\{\mathcal{E}_{pos}\cup\mathcal{E}_{neg}\}$, is the set of evaluation edges, for example the test edges, which comprises both positive edges $\mathcal{E}_{pos}$, and negative edges $\mathcal{E}_{neg}$. Roughly speaking, $\nabla$ contains, the edge gradients of positive and negative test edges computed at each layer of the model. This is done only for message-passing layers, excluding those in the Encoding and Decoding phase, in order to isolate the effect of the message-passing. We specify the squared norm of the positive (or negative) edge gradients at time t as follows:

### v3

@@ Abstract
The message-passing mechanism underlying Graph Neural Networks (GNNs) is not naturally suited for heterophilic datasets, where adjacent nodes often have different labels. Most solutions to this problem remain confined to the task of node classification. In this article, we focus on the valuable task of link prediction under heterophily, an interesting problem for recommendation systems, social network analysis, and other applications. GNNs like GRAFF have improved node classification under heterophily by incorporating physics biases in the architecture. Similarly, we propose GRAFF-LP, an extension of GRAFF for link prediction. We show that GRAFF-LP effectively discriminates existing from non-existing edges by learning implicitly to separate the edge gradients. Based on this information, we propose a new readout function inspired by physics. Remarkably, this new function not only enhances the performance of GRAFF-LP but also improves that of other baseline models, leading us to reconsider how every link prediction experiment has been conducted so far. Finally, we provide evidence that even simple GNNs did not experience greater difficulty in predicting heterophilic links compared to homophilic ones. This leads us to believe in the necessity for heterophily measures specifically tailored for link prediction, distinct from those used in node classification. The code and appendix are available at https://github.com/difra100/Link_Prediction_with_PIGNN_IJCNN.
11footnotetext: Department of Computer Science, Control and Management Engineering, Sapienza University of Rome, Rome, Italy.22footnotetext: Institute of Information Science and Technologies ”Alessandro Faedo” - ISTI-CNR, Pisa, Italy.§§footnotetext: Corresponding author, difrancesco@diag.uniroma1.it.
@@ I Introduction
Most GNNs rely on the message-passing formalism [1]. In heterophilic graphs, where connected nodes tend to have different labels, Message-Passing Neural Networks (MPNNs) may struggle in the classification task as they tend to generate similar representations for adjacent nodes [2], an issue commonly known as over-smoothing [3]. There have been efforts to improve the performance of GNNs on heterophilic graphs [4], however, these works are limited to the context of node classification. An example of this is the gradient flow framework (GRAFF) [5], which deals with heterophily through physics-inspired biases. Although GRAFF shows competitive performance in both heterophilic and homophilic graphs, existing research has focused solely on node classification tasks, not investigating its potential in link prediction, which has significant applications in several domains. This task becomes particularly complex for heterophilic graphs, where we may have a link between two nodes with dissimilar characteristics. Unlike homophilic graphs, where connected nodes are similar, the reason behind the connection of two dissimilar entities can be latent. To the best of our knowledge, link prediction under heterophily has only been discussed and brought to the community’s attention by [6].
@@
    We propose a novel Physics-Inspired readout function that leads to consistent performance improvements for GRAFF-LP as well as other baseline models. Additionally, the new readout gives GRAFF-LP more transparency in its behavior at inference time.
@@ II Related Works
Physics-Inspired vs. Physics-Informed. Physics-Informed (PI) neural networks incorporate physical priors to improve performance and generalizability [12]. Physics-Inspired (PIrd) networks are a subset of PI methods, where physical constraints are embedded in the architecture itself, acting as inductive biases [13].
Physics-Inspired Graph Neural Networks. PIrd GNNs incorporate physics principles directly into the model’s structure. Examples include GNNs based on gradient flows [5], reaction-diffusion equations [14], based on nonlinear controlled and damped oscillators [13]. While these approaches have been applied to node classification, they remain unexplored for link prediction. Our work provides the first perspective on PIrd biases in this setting. The extended discussion on related works is available in Appendix Extended Related Works.
@@ III Preliminaries
Homophily measures. The homophily assumption in graphs refers to the tendency of similar nodes to be connected. An unambiguous similarity measure is missing in the literature. Those that have been used the most with GNNs are edge homophily $\xi_{edge}$ [15] and node homophily $\xi_{node}$ [2]. The former was more considered within the node classification benchmarks [4, 5], and can be computed as
@@
Both $\xi_{edge}$ and $\xi_{node}$ rely on the labels associated with the nodes. As an example, Equation (1) measures the fraction of edges that connect nodes from the same class. Generally, if we record a low homophily, we consider the graph as heterophilic. Traditional homophily measures are unsuitable for cross-datasets comparison because they are sensitive to class numbers and sample balance [16, 7]. To address this, the adjusted homophily metric $\xi_{adj}$ was introduced in [16]:
@@
Graph Neural Networks as gradient flow. Let us consider an $N$-dimensional dynamic system evolving as $\dot{\mathbf{H}}(t)=F(\mathbf{H}(t))$, with $\mathbf{H}(t)\in\mathbb{R}^{N\times d}$ If there exists a function $E:\mathbb{R}^{N\times d}\rightarrow\mathbb{R}$, s.t. $F(\mathbf{H}(t))=-\nabla E(\mathbf{H}(t))$, the evolution equation $\dot{\mathbf{H}}(t)$ is the gradient flow of the energy $E$. Gradient flows are useful for studying the underlying dynamics of the system, provided the knowledge of $E$. In our case, $N$ represents the graph nodes whose representations $\mathbf{H}(t)$ evolve through a GNN over time, and $E(\mathbf{H}(t))$, is an energy functional associated with the node representations. Let $GNN:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ be an intermediate layer of a generic GNN. By treating the GNN layers as continuous time $t$ and defining $GNN(\mathbf{H}(t))=-\nabla E(\mathbf{H}(t))$, the evolution of the features through the GNN is described as the gradient flow of $E(\mathbf{H}(t))$. $E$ can be selected as the Dirichlet Energy $E^{dir}$:
@@
where $(\nabla\mathbf{H}(t))_{ij}=\frac{\mathbf{h}^{t}_{j}}{\sqrt{D_{jj}+1}}-\frac{\mathbf{h}^{t}_{i}}{\sqrt{D_{ii}+1}}$ and $\mathbf{h}^{t}_{i}$ denotes the feature of node $i$ at time $t$. Poor performance on heterophilic graphs and over-smoothing are often linked to the Dirichlet energy of features decaying to zero as layers increase [17, 18]. Equation (3) shows that decreasing $E^{dir}$ brings adjacent nodes closer in feature space. Conversely, from what was commonly thought, [5] proved that linear graph convolutions with symmetric weights shared among layers can induce edge-wise attraction (repulsion) through their positive (negative) eigenvalues. This control mechanism effectively influences whether the features are smoothed or sharpened, making the model successfully handle node classification within heterophilic graphs. This was possible by defining layers as a gradient flow of a parametrized Dirichlet energy:
@@
$\Theta_{+}$ and $\Theta_{-}$ depend on the positive and negative eigenvalues of the weight matrices, respectively. By applying Euler discretization to the gradient flow and replacing $\mathbf{H}(t)$ by $\mathbf{H}^{t}$:
@@
$\tau$ is the integration step, $T=\tau L$ is the total integration time, $\mathbf{\Omega}$, $\mathbf{W}$ and $\tilde{\mathbf{W}}$ are the trainable matrices, that are symmetric and shared across the layers, and $\boldsymbol{\mathbb{A}}=\mathbf{\tilde{D}}^{-\frac{1}{2}}\mathbf{\tilde{A}}\mathbf{\tilde{D}}^{-\frac{1}{2}}$, since self-loops are included in $\mathbf{A}$ and $\mathbf{D}$ (i.e. $\tilde{\mathbf{A}}=\mathbf{I}+\mathbf{A}$). Equation (5) takes the form of a residual network [19], where the interpretation of GNNs as gradient flows is referred to in the literature as PIrd GNN. For details on these derivations, see [5]. In this work, we take advantage of these results and build upon this architecture for link prediction.
@@ IV Proposed Framework: GRAFF-LP
The Hadamard readout is commonly used with node-based link predictors [8]. In this paper, we are the first to propose an alternative function $f_{g}$, which we believe is more compatible with the GRAFF backbone. As seen in Equation (4), gradient flow operates by minimizing or maximizing the squared norm of edge gradients, and $f_{g}$ can be directly related to them through the following relation:
@@
More precisely, the gradient flow operates by minimizing or maximizing the squared norm of the gradients multiplied by the terms $\Theta_{+}$ and $\Theta_{-}$; we preferred to define the readout function without including those terms to speed up the calculation of $f_{g}$. We use the Hadamard product of the gradients to avoid any bias on the edge directionality since we deal with undirected graphs. However, for directed graphs, we could simply have $f_{g}(\mathbf{z}_{i},\mathbf{z}_{j})=(\nabla\mathbf{H}^{T})_{i,j}$. We do not consider other readouts such as concatenation of $z_{i}$ and $z_{j}$, since this would provide a bias on directionality as well, we would lose the physics-inspired bias, and we would double $d_{MLP}$, negatively affecting the space and time complexity.
@@ IV-A Complexity Analysis of GRAFF-LP
We now consider the space complexity in terms of the number of parameters. GRAFF-LP, due to its weight-sharing mechanism, maintains constant complexity w.r.t. $L$ and exhibits quadratic complexity w.r.t. the number of hidden dimensions, as the elements in $\mathbf{\Omega}$ and $\mathbf{W}$ are of size $d_{h}^{2}$, where $d_{h}$ is the hidden dimension. Since these matrices are symmetric, there is redundancy in the parameters, reducing the total number to approximately $\frac{1}{2}d_{h}^{2}$.
To summarize, the number of parameters for GCN scales as $\mathcal{O}(Ld_{h}^{2})$. Since GCN shares a similar message-passing as Equation (5), but it does not use weight sharing, and we set $\mathbf{\Omega}\equiv 0$. In the case of GRAFF-LP, both $\mathbf{W}$ and $\mathbf{\Omega}$ use weight sharing, and the overall complexity scales as $\mathcal{O}(d_{h}^{2})$. In terms of time complexity, GRAFF-LP does not have any specific advantage over other models. In Table V, we show the runtime analysis for all the models, reporting also the number of parameters.
@@ V-A Experimental Set-up
V-A1 Datasets
Table I presents graph statistics after conversion to undirected graphs, a standard GNN procedure. The datasets selected belong to a recent collection [7], which was proposed to enrich the current dataset availability for the GNN experimental setting under heterophily. They have not yet been applied to link prediction. We used four datasets, Amazon Ratings, Roman Empire, Minesweeper, Questions, and Tolokers. Additional details and dataset descriptions can be found in Appendix Datasets Description. These datasets differ in context, size, and structural properties [7].
@@ V-A2 Baselines
GAT: An attentional MPNN [21] that uses an attention mechanism for neighbor aggregation, particularly effective for heterophilic graphs.
@@
NCNC: [23] A link prediction method, based on structural features as ELPH, but these are related to higher-order common neighbor information. This method is currently state-of-the-art in a recent link prediction benchmark [11].
@@
Further experimental details are available in Appendix Experimental Set-up: Additional Details.
@@ V-B1 Effectiveness
To evaluate model performance, we use the Area Under the Receiver Operating Characteristic (AUROC). Table II summarizes the results across all datasets for both readout functions, $f_{h}$ and $f_{g}$. The results with $f_{h}$ are related to the best configuration of hyperparameters with $f_{h}$. Concerning $f_{g}$, we just took the same configuration found with $f_{h}$ and trained the models again via $f_{g}$. Performance metrics are averaged over 10 random seeds. We adopted a single split since the graphs are sufficiently large and less sensitive to high variance in the data. The statistical significance of the results is computed through the Wilcoxon test [24], via the accuracy of positive and negative edges in the test set; details about the procedure can be found in the code.
@@
With $f_{h}$, GRAFF-LP leads in 3 out of 4 datasets and in 2 out of 4 using $f_{g}$, ranking consistently in the top 2 across all datasets, indicating adaptability not seen with the other models. These aspects also apply to Tolokers, whose results and comments are reported in Appendix Additional Results.
@@
On the Roman Empire dataset, which resembles a chain-like graph, GRAFF-LP achieves the highest AUROC in both setups, with a neat advantage w.r.t. other baselines. ELPH, and NCNC also present a wide gap w.r.t node-based methods, but they have access to the subgraph features that are a clear advantage to perform link prediction in such a structure. This advantage and superiority of ELPH and NCNC are found with Minesweeper, where the graph is a regular grid, and link prediction can be easily solved by understanding the grid structure.
@@
In both cases, the physics-inspired bias seems to let the model use the structure properly and achieve state-of-the-art performance. According to our hyperparameter optimization, ELPH fails to obtain competitive performance in Amazon Ratings and Questions. We relate this gap to the limited depth of ELPH (which is up to 3 layers). NCNC instead can reach competitive performance in Amazon Ratings, as well as Questions. Despite this, it never surpasses GRAFF-LP. We conclude that subgraph-based methods do not perform better than GRAFF-LP on class heterophilic datasets. This implies that structural features are not critical or determining to perform well in these settings.
@@
Another observation is that GCN, both in Amazon Ratings and Questions, is among the top-scoring models, and in particular gets the closest to GRAFF-LP. This is not unexpected, since they share a similar message-passing and expressivity [25, 5]. They mainly differ in the weight design. We associate these similarities to the same phenomenon observed by [5], who observed that in homophilic node classification, the two models perform the same. We conjecture in this case that for Amazon Ratings and Questions the edge-wise repulsion is not required, thus GCN can afford similar results. We also report the results of the MLP. The MLP results consistently lag behind the GNN models, emphasizing the need for graph representation learning for effective link prediction in these new datasets. For example, in Minesweeper, MLP struggles to infer missing links due to its lack of graph context and the dataset’s grid structure.
@@
In Appendix -D, we provide examples of GRAFF-LP performing on link prediction under homophily, to show that our approach can reach competitive as well as state-of-the-art performance, also under homophily.
In Table III, we show the relative percentage improvement we obtained using $f_{g}$ as the readout method. We did not include ELPH and NCNC, since their official derivation is based on Hadamard or other readouts.
@@
GRAFF-LP +0.70% +1.02% 0% 0%
V-B2 Physics-Inspired Link Prediction
[5] proposed the GRAFF architecture and showed the attraction and repulsion behavior among adjacent nodes with different labels. This is expressed in Equation (4). In the link prediction scenario, we believe that the same behavior can be induced among two nodes that, even though they seem different, should attract themselves, and vice versa, where two nodes are similar but do not present any interaction. Since we build upon the same framework of [5], we expect the edge gradients associated with $\Theta_{+}$ to become smaller, because of attraction in the feature space, and the edge gradients associated with $\Theta_{-}$ to become larger in the feature space because of repulsion. In this way, we can deduce that GRAFF-LP can induce attraction and repulsion in the same fashion as in GRAFF for node classification, and we would answer positively to RQ2. To understand whether this phenomenon happens, we introduce a novel metric, namely the gradient separability ($GS$). Let us assume we have a set of edge gradients =$\nabla=\{(\nabla\mathbf{H}^{t})_{i,j}|0\leq t\leq T\wedge(i,j)\in\mathcal{E}_{t}\}$, where $\mathcal{E}_{t}=\{\mathcal{E}_{pos}\cup\mathcal{E}_{neg}\}$, is the set of evaluation edges, for example the test edges, which comprises both positive edges $\mathcal{E}_{pos}$, and negative edges $\mathcal{E}_{neg}$. Roughly speaking, $\nabla$ contains the edge gradients of positive and negative test edges computed at each layer of the model. This is done only for message-passing layers, excluding those in the Encoding and Decoding phases, to isolate the effect of the message-passing. We specify the squared norm of the positive (or negative) edge gradients at time t as follows: